# CBMC-V3: A CNS-inspired Control Framework Towards Manipulation Agility with SNN

### v2

@@ II-B Hierarchical control level
Clearly, different neural regions in the central nervous system operate hierarchically during motion control, with each level specializing in particular tasks while communicating and cooperating through neural pathways. This hierarchical organization enables organisms to execute a wide range of motion tasks flexibly and stably across varying environments. Physiological studies indicate that the multilevel transmission of neural signals is a key component of this hierarchy. Proprioceptive information is conveyed from sensory receptors to the brain via a three-order neuronal pathway[william_ascending_pathway, LLOBERA2023190], as illustrated in Fig.1a. First-order neurons receive impulses from the skin and muscles and transmit them to the spinal cord. Some of this information is processed locally in the spinal cord and then relayed back to the muscles through these first-order neurons. The remaining signals pass through second-order neurons, which transmit impulses via the brainstem to the thalamus. Finally, third-order neurons carry these impulses to the cerebral cortex and cerebellum, and are responsible for information exchange between them.
@@
  • First-order level: Comprising the spinal cord module, this level operates at the highest frequency. Its primary function is to generate real-time control torques sent to the robotic arm, ensuring accurate and rapid responses.
@@
  • Second-order level: Including the brainstem and thalamus module, this level operates at an intermediate frequency. It serves as a relay between the first-order and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order level and integrating and filtering the torques provided by the third-order level to supply feedforward torque to the first-order level.
@@
  • Third-order level: Comprising the cerebellum and cerebral cortex module, this layer operates at the lowest frequency. Its role is to generate movement trajectories and provide gravity compensation torques based on motion patterns, maintaining movement stability and smoothness.
II-C Information pathway
@@
As illustrated in Fig.1a, the human central nervous system employs ascending and descending pathways for sensory information transmission and motor command execution, respectively. The ascending pathways primarily convey peripheral sensory information to the brain for perception and integration, while the descending pathways transmit motor commands from the brain to the spinal cord and peripheral nerves to control movements. Following this principle, our framework incorporates two analogous information pathways:
@@
  • Ascending pathway: Inspired by the dorsal column–medial lemniscal and spinocerebellar tracts[william_ascending_pathway], this pathway transmits proprioceptive and touch information, namely joint positions, velocities, and end-effector torque. It simulates the process by which muscle receptors relay signals to the CNS for processing and can be regarded as the input process of the framework.
@@
  • Descending pathway: Inspired by the corticospinal and reticulospinal tracts[william_ascending_pathway], this pathway transmits joint torque commands and parameter adjustment signals. It simulates the process by which the CNS delivers motor commands to muscle effectors and can be regarded as the output process of the framework.
II-D A complete operational cycle
To clarify the interplay among modules, levels, and pathways, we illustrate in Fig. 1b a complete working cycle of the proposed framework, showing how sensory inputs are transformed into motor control outputs.
In each cycle, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$ and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$) are transmitted via first-order neurons to the spinal cord. Part of this input is locally processed to generate feedback torque $\bm{\tau}_{\text{fb}}$; part is relayed via second-order neurons to the brainstem, which adaptively provides PD gains $(\bm{K}_{P},\bm{K}_{V})$; and part is sent via third-order neurons to the cerebellum, which outputs gravity compensation matrix $\bm{T}_{\text{cb}}$. These ascending transmissions complete the flow of proprioceptive information from sensors to the modules.
On the descending side, $\bm{T}_{\text{cb}}$ is conveyed via third-order neurons to the thalamus, where end-effector torque feedback $\bm{F}_{\text{ef}}$ is integrated to generate a weight vector $\bm{w}$. This vector filters $\bm{T}_{\text{cb}}$ to produce the effective gravity compensation torque $\bm{\tau}_{g}$, which is passed via second-order neurons to the spinal cord. There, $\bm{\tau}_{g}$ is combined with $\bm{\tau}_{\text{fb}}$ to form the final control torque $\bm{\tau}$, which is delivered through first-order neurons to the actuators. Thus, the descending pathway completes motor command transmission from the modules to the joints.
Because information reaches different modules through pathways of varying length, the framework is naturally organized into three control levels. The first-order level involves only spinal processing and yields the shortest latency and highest frequency. The second-order level adds brainstem and thalamic processing, resulting in longer pathways and lower frequency. The third-order level incorporates the cerebellum, producing the longest transmission path and the lowest frequency.
@@ II-D A complete operational cycle
The pseudocode implementation of this framework is provided in Algorithm 1.
@@
0:  joint proprioceptive sensor data $\bm{\dot{q}},\bm{q}$, planned trajectory data $\bm{\dot{q}}_{d},\bm{q}_{d}$, end-effector torque sensor data $\bm{F}_{ef}$
@@
6:  initialize intermediate variables $\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g},\bm{w},\bm{T}_{{\text{cb}}}$
7:  read trajectory length $M$
8:  for $t=1$ to $M$ do
@@
22:
23:  Sub-thread 1:
24:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q},\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g}$
25:  process through the spinal cord module
26:  update $\bm{\tau}$
27:
28:  Sub-thread 2:
29:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
30:  process through the brainstem module
31:  update $\bm{K}_{V},\bm{K}_{P}$
32:
33:  Sub-thread 3:
34:  read $\bm{T}_{{\text{cb}}},\bm{F}_{ef}$
35:  process through the thalamus module
36:  update $\bm{\tau}_{g},\bm{w}$
37:
38:  Sub-thread 4:
39:  read $\bm{q}$
40:  process through the cerebellum module
41:  update $\bm{T}_{{\text{cb}}}$
III Module network design
Building on the framework introduced in the previous section, this chapter focuses on the implementation of each module using SNNs. The cerebral cortex module, which serves only for trajectory storage and playback, is therefore excluded from the discussion. Since high-performance robotic arms are commonly designed with seven DoF, all implementations in this chapter target a 7-DoF arm. For manipulators with different DoF, the same design principles can be applied by appropriately scaling the neural network architecture of each module.
@@ III-A Spiking neuron
SNN employs spiking neurons as their fundamental computational units. A variety of spiking neuron models have been proposed in the literature, such as the Hodgkin–Huxley model [hodgkin1952quantitative] and the Leaky Integrate-and-Fire (LIF) model along with its derivatives [burkitt2006review]. The LIF model is of particular interest here, since it not only captures the dynamic process of membrane potential accumulation and leakage, but also maintains relatively high computational efficiency [stein1965theoretical]. Therefore, we will focus on the LIF model.
@@
$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)\exp\left(-s/\tau_{s}\right)\text{d}s}$ (7)
where $\tau_{s}$ denotes the synaptic time constant. The membrane potential $u$ of the neuron evolves according to the following dynamics:
$\tau_{m}\frac{\text{d}u}{\text{d}t}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
where $\tau_{m}=RC$ is the membrane time constant determined by the membrane resistance $R$ and capacitance $C$. Here, $u_{\text{reset}}$ is the reset potential after firing, $i_{0}(t)$ represents the external current driving the neuron, $i_{j}(t)$ denotes the input current from the $j$-th synapse, and $w_{j}$ is the synaptic weight associated with the $j$-th synapse. When the membrane potential $u$ reaches a certain threshold $u_{\text{fire}}$, the neuron emits a spike and the potential is reset to $u_{\text{reset}}$. Therefore, the operational process of a spiking neuron can be interpreted as follows:
1.When the membrane potential $u(t)$ exceeds the firing threshold $u_{\text{fire}}$, the neuron emits a spike $s(t)$, expressed as
@@
2.At each time step, the differential equation (8) is updated, where the membrane potential decays by $\lvert u_{\text{reset}}-u(t)\rvert$ (representing the forgetting effect over time) and increases by $R\left(i_{0}(t)+\sum w_{j}i_{j}(t)\right)$ (representing the stimulation effect induced by external inputs).
3.The process then returns to Step 1.
In our framework, we adopt a backpropagation-based training method[8891809] for SNN. Since the activation function of the neuron in equation (10) is a discontinuous function, directly computing its derivative can lead to network instability. Consequently, various surrogate gradient methods have been proposed [8891809]. A common approach is to use the original step function $y=\Theta(a)$ during the forward pass, while replacing its derivative with $\theta^{\prime}(a)=\sigma^{\prime}(a)$ instead of $\Theta^{\prime}(a)$ during backpropagation, where $\sigma(a)$ is referred to as the surrogate activation function. Typically, $\sigma(a)$ is a smooth and continuous function that resembles the shape of $\Theta(a)$ while enabling stable gradient computation.
As illustrated in Fig.1c, the framework employs three types of neurons across all modules, represented in gray, yellow, and green. The gray neurons correspond to input fibers, which serve as the data interface for each module via connections to the ascending or descending pathway. The yellow neurons represent LIF neurons, which constitute the core units of SNNs and are responsible for information storage, processing, and learning. The green neurons correspond to non-spiking LIF neurons, implemented by setting the membrane time constant $\tau_{m}$ of standard LIF neurons to infinity. Unlike spiking neurons, their output is the continuous membrane potential $u(t)$ rather than discrete spikes, thereby enabling spike-to-continuous signal conversion. These neurons serve as the output layer of each module, connecting to the descending pathway, and, in addition to the functional capabilities of the yellow LIF neurons, they also perform information decoding and output.
@@ III-B Cerebellum module
Humans are capable of rapidly acquiring complex skills through imitation, and neural networks exhibit similar properties. Therefore, compared with unsupervised learning approaches, training a neural network to perform gravity compensation via imitation provides a faster and more efficient solution. Under the condition of a known robotic arm dynamics model, the gravity compensation torque represents a nonlinear mapping from joint positions to joint torques. In the machine learning domain, the task of fitting a curve to observed data under supervision is referred to as a regression task, for which extensive research and mature methods already exist[henkes2024spiking, robotics13090126].
Figure 2: (a) Structure of one network unit in the cerebellum module during training. (b) Structure of cerebellum network during prediction.
@@
As shown in Fig.2a, to implement the cerebellum module, a four-layer recurrent SNN is designed as a network unit of the module, analogous to the “microcomplex” proposed in [abadia2019robot, abadia2021cerebellar]. The input variable $\bm{q}\in\mathbb{R}^{7}$ corresponds to the actual joint positions of the robotic arm, and the output $\bm{\tau}_{\text{cb}}\in\mathbb{R}^{7}$ represents the gravity compensation torques under a given load condition.
@@
Variables from the ascending pathway first arrive at the mossy fiber (MF) layer, which is fully connected to the granule cell (GC) layer consisting of 300 LIF neurons. The GC layer, in turn, is fully connected to the Purkinje cell (PC) layer containing 200 LIF neurons. Each PC neuron receives input from both the current and previous time step of the GC layer, capturing temporal dynamics. The output of the PC layer is fully connected to the deep cerebellar nuclei (DCN) layer, which consists of 140 non-spiking LIF neurons. The decoding of the DCN layer is inspired by population coding observed in the motor cortex of animals [georgopoulos1986neuronal]. Studies have shown that during arm movement in primates, the activity of any single motor cortex neuron does not directly indicate the actual movement direction; however, a weighted sum of vectors corresponding to the preferred directions of a population of neurons yields a vector aligned with the movement direction. Analogously, the 140 DCN neurons are grouped in sets of 20, and their membrane potentials are averaged to obtain a 7-dimensional torque output $\bm{\tau}_{\text{cb}}$, i.e.,
$\tau_{\text{cb},i}=\sum_{j=20\cdot(i-1)}^{20\cdot i}u_{\text{DCN},j}$ (11)
@@
where $\tau_{\text{cb},i},i=1,\cdots,7$ represents the torque for each joint and $u_{\text{DCN},j},j=1,\cdots,140$ denotes the membrane potential of each DCN neuron. The network employs rate-based encoding[gerstner2002spiking], with a time window of length $T$.
@@
The training data $\bm{\tau}_{\text{ref}}$ are generated using the robotic arm’s dynamics model. Under the condition that the training data adequately cover the workspace, this regression-based approach is sufficient to achieve effective gravity compensation. The network is trained in a supervised manner, with the objective of minimizing the error between the torques output by the network and the reference torques computed from the dynamics model. Structurally, this is reflected in the DCN layer neurons receiving both excitatory inputs from the PC layer neurons and inhibitory inputs from the climbing fiber (CF) neurons. Mathematically, the network output torque $\bm{\tau}_{\text{cb}}$ is compared with the reference torque $\bm{\tau}_{\text{ref}}$ to obtain an error, which is quantified using the mean squared error (MSE Loss), i.e.,
@@
The weights are then updated via backpropagation, thereby completing the training process; detailed procedures are provided in Algorithm 2.
Algorithm 2 Training process of one network unit in the cerebellum module
@@
Training parameter Value Neuron parameter Value
@@
After training, the network can operate without reference torques, with CF neurons no longer required, relying solely on the learned synaptic weights to generate gravity compensation torques. To accommodate different loads, two network units are trained separately under light (no load) and heavy (3 kg load) modes using the supervised learning procedure described in Algorithm 2. These trained networks are then combined to form the complete cerebellum module structure illustrated in Fig. 2b.
In summary, after training, the cerebellum module receives as input the actual joint positions of the robotic arm, $\bm{q}\in\mathbb{R}^{7}$, and outputs a gravity compensation matrix composed of two torque vectors, i.e., $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$, $\bm{\tau}_{\text{cb,L}}\in\mathbb{R}^{7}$ and $\bm{\tau}_{\text{cb,H}}\in\mathbb{R}^{7}$. This matrix corresponds to the learned patterns of the cerebellum module and provides gravity compensation for trajectory tracking tasks under both light and heavy conditions. The outputs are subsequently integrated by the thalamus module to achieve gravity compensation for arbitrary loads within a specified range. During trajectory tracking, the prediction process of the cerebellum module is described in Algorithm 3, which can be implemented as Sub-thread 4 in Algorithm 1.
@@
1:  load the network for light and heavy load conditions
@@
4:  feed $\bm{q}$ into the two trained network unit, and generate two output $\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}}$

### v3

@@ II-B Control levels
Physiological studies indicate that the multilevel transmission of neural signals is a key component of this hierarchy. Proprioceptive information is conveyed from sensory receptors to the brain via a three-order neuronal pathway[28, 30, 40], as illustrated in Fig.1a. First-order neurons receive impulses from the skin and muscles and transmit them to the spinal cord. Some of this information is processed locally in the spinal cord and then relayed back to the muscles through these first-order neurons. The remaining signals pass through second-order neurons, which transmit impulses via the brainstem to the thalamus. Finally, third-order neurons carry these impulses to the cerebral cortex and cerebellum, and are responsible for information exchange between them.
@@
  • First-order level: Comprising the spinal cord module, this level operates at the highest frequency. It generates real-time control torques sent to the robotic arm, ensuring accurate and rapid responses.
@@
  • Second-order level: Including the brainstem module and thalamus module, this level operates at an intermediate frequency. It serves as a relay between the first-order and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order level, and integrating and filtering the torques provided by the third-order level to supply gravity compensation torque to the first-order level.
@@
  • Third-order level: Comprising the cerebellum module and cerebral cortex module, this layer operates at the lowest frequency. It generates motion trajectories and provide compensation torques based on motor patterns, maintaining movement stability and smoothness.
II-C Information pathways
@@
As illustrated in Fig.1a, the human CNS employs ascending and descending pathways for sensory information transmission and motor command execution, respectively[40, 30]. The ascending pathway conveys peripheral sensory information to the brain for perception and integration, while the descending pathway transmits motor commands from the brain to the spinal cord and peripheral nerves to control movements. Following this principle, our framework incorporates two analogous information pathways:
@@
  • Ascending pathway: This pathway transmits proprioceptive and touch information, namely joint positions, velocities, and end-effector torque feedback. It simulates the process by which muscle receptors relay signals to the CNS for processing and can be regarded as the input process of the framework.
@@
  • Descending pathway: This pathway transmits joint torque commands and module coefficients. It simulates the process by which the CNS delivers motor commands to muscle effectors and can be regarded as the output process of the framework.
II-D Pseudocode implementation
@@
To clarify the interplay among modules, levels, and pathways illustrated in Fig. 1b, the pseudocode implementation of this framework is provided in Algorithm 1.
Information travels different length to reach different modules along the pathway, so the framework is organized into three control levels with different frequencies.
On the ascending side, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$ and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$) are transmitted via first-order neurons to the spinal cord module. Part of this input is locally processed to generate feedback torque $\bm{\tau}_{\text{fb}}$; part is relayed via second-order neurons to the brainstem module, which adaptively provides feedback coefficients $(\bm{K}_{P},\bm{K}_{V})$; and part is sent via third-order neurons to the cerebellum module, which outputs compensation matrix $\bm{T}_{\text{cb}}$.
On the descending side, $\bm{T}_{\text{cb}}$ is conveyed via third-order neurons to the thalamus module. It uses end-effector torque feedback $\bm{F}_{\text{ef}}$ to generate a weight vector $\bm{w}$ which filters $\bm{T}_{\text{cb}}$ to produce the gravity compensation torque $\bm{\tau}_{g}$. This torque is passed via second-order neurons to the spinal cord module. In the spinal cord module, $\bm{\tau}_{g}$ is combined with $\bm{\tau}_{\text{fb}}$ to form the control torque command $\bm{\tau}$, which is delivered through first-order neurons to the actuators.
@@
0:  actual joint velocity and position $\bm{\dot{q}},\bm{q}$, planned joint velocity and position $\bm{\dot{q}}_{d},\bm{q}_{d}$, end-effector torque sensor data $\bm{F}_{ef}$
@@
6:  initialize intermediate variables $\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g},\bm{T}_{{\text{cb}}}$
7:  read trajectory length $L$
8:  for $t=1$ to $L$ do
@@
22:  Sub-thread 1:
23:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q},\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g}$
24:  process through the spinal cord module
25:  update $\bm{\tau}$
26:  Sub-thread 2:
27:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
28:  process through the brainstem module
29:  update $\bm{K}_{V},\bm{K}_{P}$
30:  Sub-thread 3:
31:  read $\bm{T}_{{\text{cb}}},\bm{F}_{ef}$
32:  process through the thalamus module
33:  update $\bm{\tau}_{g}$
34:  Sub-thread 4:
35:  read $\bm{q}$
36:  process through the cerebellum module
37:  update $\bm{T}_{{\text{cb}}}$
III Module design
Building on the framework introduced in the previous section, this chapter focuses on the implementation of each module using SNNs. The cerebral cortex module, which serves only for trajectory storage and playback, is therefore excluded from the discussion. All implementations in this chapter target a 7-DOF arm, i.e., $n=7$.
@@ III-A Spiking neuron
SNN employs spiking neurons as its computation units. A variety of spiking neuron models have been proposed in the literature, such as the Hodgkin–Huxley model [22] and the Leaky Integrate-and-Fire (LIF) model along with its derivatives [7]. The LIF model is of particular interest here, since it not only captures the dynamic process of membrane potential accumulation and leakage, but also maintains relatively high computational efficiency [41].
@@
$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)\exp\left(-s/\tau_{\text{s}}\right)\text{d}s}$ (7)
where $\tau_{\text{s}}$ denotes the synaptic time constant. The membrane potential $u(t)$ of the neuron evolves according to the following dynamics:
$\tau_{\text{mb}}\frac{\text{d}u}{\text{d}t}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
where $\tau_{\text{mb}}=RC$ is the membrane time constant determined by the membrane resistance $R$ and capacitance $C$. Here, $u_{\text{reset}}$ is the reset potential after firing, $i_{0}(t)$ represents the external current driving the neuron, $i_{j}(t)$ denotes the input current from the $j$-th synapse, and $w_{j}$ is the synaptic weight associated with the $j$-th synapse. When the membrane potential $u$ reaches a certain threshold $u_{\text{fire}}$, the neuron emits a spike and the potential is reset to $u_{\text{reset}}$. Therefore, the operational process of a spiking neuron can be interpreted as follows:
1. When the membrane potential $u(t)$ exceeds the firing threshold $u_{\text{fire}}$, the neuron emits a spike $s(t)$, expressed as
@@
2. Update equation (8).
3. Return to Step 1.
In our framework, we adopt a back-propagation-based training method[31] for SNN. Since the activation function of the neuron in equation (10) is a discontinuous function, directly computing its derivative can lead to network instability. Consequently, various surrogate gradient methods have been proposed [31]. A common approach is to use the original step function $\Theta(x)$ during the forward pass, while replacing its derivative with $\sigma^{\prime}(x)$ instead of $\Theta^{\prime}(x)$ during back propagation, where $\sigma(x)$ is referred to as the surrogate activation function. Typically, $\sigma(x)$ is a smooth and continuous function that resembles the shape of $\Theta(x)$ while enabling stable gradient computation.
As illustrated in Fig.1c, the framework employs three types of neurons across all modules, represented in gray, yellow, and green. The gray neurons correspond to input fibers, which serve as the data interface for each module via connections to the ascending or descending pathway. The yellow neurons represent LIF neurons, which constitute the core units of SNNs and are responsible for information storage, processing, and learning. The green neurons correspond to non-spiking LIF neurons, implemented by setting the firing threshold $u_{\text{fire}}$ of standard LIF neurons to infinity. Unlike spiking neurons, their output is the continuous membrane potential $u(t)$ rather than discrete spikes, thereby enabling spike-to-continuous signal conversion. These neurons, connecting to the descending pathway, serve as the decoding layer of each module.
@@ III-B Cerebellum module
Humans are capable of rapidly acquiring complex skills through imitation, and neural networks exhibit similar properties. Therefore, compared with unsupervised learning approaches, training a neural network to perform gravity compensation via imitation provides a faster and more efficient solution. Under the condition of a known robotic arm dynamic model, the gravity compensation torque represents a nonlinear mapping from joint positions to joint torques. In the machine learning domain, the task of fitting a curve to observed data under supervision is referred to as a regression task, for which extensive research and mature methods already exist[21, 42].
@@
As shown in Fig.2a, to implement the cerebellum module, a four-layer recurrent SNN inspired by [2, 3] is designed to store one motor pattern corresponding to a particular load mass. The input variable $\bm{q}\in\mathbb{R}^{7}$ corresponds to the actual joint positions of the robotic arm, and the output $\bm{\tau}_{\text{cb}}\in\mathbb{R}^{7}$ represents the gravity compensation torque under a given load condition.
Figure 2: Structure of the cerebellum module. (a) In training. (b) In prediction.
@@
Variables from the ascending pathway first arrive at the mossy fiber (MF) layer, which is fully connected to the granule cell (GC) layer consisting of 300 LIF neurons. The GC layer, in turn, is fully connected to the Purkinje cell (PC) layer containing 200 LIF neurons. Each PC neuron receives input from both the current and previous time step of the GC layer, capturing temporal dynamics. The output of the PC layer is fully connected to the deep cerebellar nuclei (DCN) layer, which consists of 140 non-spiking LIF neurons. The decoding of the DCN layer is inspired by population coding in the animal motor cortex [16]. During primate arm movements, individual motor cortex neurons do not directly represent movement direction; rather, the movement direction emerges from a weighted vector sum over the preferred directions of a neuronal population. Analogously, the 140 DCN neurons are grouped in sets of 20, and their membrane potentials are averaged to obtain a 7-dimensional torque output $\bm{\tau}_{\text{cb}}$, i.e.,
$\tau_{\text{cb},i}=\sum_{j=20\cdot(i-1)+1}^{20\cdot i}u_{\text{DCN},j}$ (11)
@@
where $\tau_{\text{cb},i},i=1,\cdots,7$ represents the torque for each joint and $u_{\text{DCN},j},j=1,\cdots,140$ denotes the membrane potential of each DCN neuron. The network employs rate-based encoding[17], with a time window of length $T$.
@@
The training data $\bm{\tau}_{\text{ref}}$ are generated using the robotic arm’s dynamic model. Under the condition that the training data adequately cover the workspace, this regression-based approach is sufficient to achieve effective gravity compensation. The network is trained in a supervised manner, with the objective of minimizing the error between the network output and the reference torque. Structurally, this is reflected in the DCN layer neurons receiving both excitatory inputs from the PC layer neurons and inhibitory inputs from the climbing fiber (CF) neurons. Mathematically, the network output torque $\bm{\tau}_{\text{cb}}$ is compared with the reference torque $\bm{\tau}_{\text{ref}}$ to obtain an error, which is quantified using the mean squared error (MSE Loss), i.e.,
@@
The weights are then updated via back propagation, thereby completing the training process. Detailed training procedure is provided in Algorithm 2. After training, the network can operate without reference torques, relying solely on the learned synaptic weights to generate gravity compensation torque.
Algorithm 2 Training process of network for one motor pattern
@@
Parameter Value Parameter Value
@@
To realize motor patterns for different load masses, two networks, as depicted in Fig. 2a, are trained separately in light pattern (with no load) and in heavy pattern (with 3 kg load) using the supervised learning procedure described in Algorithm 2. These trained networks are then combined to form the complete cerebellum module, as illustrated in Fig. 2b.
In summary, after training, the cerebellum module receives as input the actual joint positions of the robotic arm, $\bm{q}\in\mathbb{R}^{7}$, and outputs a gravity compensation matrix composed of two torque vectors, i.e., $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$, $\bm{\tau}_{\text{cb,L}}\in\mathbb{R}^{7}$ and $\bm{\tau}_{\text{cb,H}}\in\mathbb{R}^{7}$. This matrix corresponds to the two learned patterns of the cerebellum module and provides gravity compensation for trajectory tracking tasks under both light and heavy pattern. During trajectory tracking tasks, the prediction process of the cerebellum module is described in Algorithm 3, which can be implemented as Sub-thread 4 in Algorithm 1.
@@
1:  load the cerebellum module network
@@
4:  feed $\bm{q}$ into the cerebellum module network, and generate two output $\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}}$